OpenStudy (anonymous):

37.5 is 60% of what #?

OpenStudy (anonymous):

62.5

OpenStudy (anonymous):

u should explain @KamiBug do not give them websites

OpenStudy (techarena):

consider unknown number to be x. 60% of x = 37.5 (60/100) * x = 37.5 x = (37.5 * 100) / 60 = 62.5 Ref : https://www.easycalculation.com/change-value-from-perentage.php